# You Cannot Feed Two Birds with One Score: the Accuracy-Naturalness Tradeoff in Translation

### v1

@@ Appendix F Caveat with the LLM-based approximation of the AN curve
Note, that this approach will always produce an overly optimistic estimate for the curve, since we compute our estimate as
@@
where $\mathcal{L}$ is the one-shot Lagrange dual of the AN function from Equation 5, and we maximise over systems/LLM-generated candidates for each translation first, and then we average over the different entries of the test set second. However, in the true Lagrange-dual of the AN function, the order of the expectation and the maximisation is swapped: $\max\mathbb{E}[\mathcal{L}]$. Finally, we have the standard result that
@@
Appendix G Further evaluations of WMT24 shared task entries
In this section, we present a comparison of automated reference metrics versus the MQM fluency of the translations, see Appendix D for an explanation of the latter.
In particular, in Figure 4 we plot “classic” reference metrics against fluency, while in Figure 5 we plot neural metrics against fluency.
From comparing these two figures, we see that neural metrics correlate with fluency “for much longer,” that is, their accuracy-naturalness tradeoffs should be much milder. Nonetheless, our theory shows that the tradeoff exists for neural metrics too, and therefore at some point these will begin to anticorrelate with fluency as well.
Figure 4: Tradeoff of “pure” distortion metrics vs MQM fluency
Figure 5: Tradeoff of “perceptual” distortion metrics vs MQM fluency

### v3

@@ Appendix F Caveat with the LLM-based approximation of the AN curve
Note, that the approach we use to approximate the accuracy-naturalness curve in Section˜5.1 will always produce an overly optimistic estimate for the curve, since we compute our estimate as
@@
where $\mathcal{L}$ is the one-shot Lagrange dual of the AN function from Equation˜5, and we maximise over systems/LLM-generated candidates for each translation first, and then we average over the different entries of the test set second. However, in the true Lagrange-dual of the AN function, the order of the expectation and the maximisation is swapped: $\max\mathbb{E}[\mathcal{L}]$. Finally, we have the standard result that